# Generative Models in Decision Making: A Survey

### v2

@@ 4.3.7 Autoregressive Models as Optimizer
Autoregressive models, particularly due to their sequential structure, are effective optimizers because they capture temporal or sequential dependencies between data points, allowing for iterative refinement of solutions. This structure enables the models to handle uncertainty and adaptively improve the optimization performance across various tasks. The sequential nature of autoregressive models, which predicts the next step based on previous steps, allows them to generate or adjust solutions gradually, making them highly effective in settings where the solution space is complex and high-dimensional.
For instance, Transformer Neural Processes (TNPs) [185] introduce a meta-learning framework that employs a transformer-based autoregressive model for optimizing performance under uncertainty in tasks that involve variable inputs or distributions. By leveraging the strengths of transformers in sequence modeling and combining them with Neural Processes for better uncertainty estimation, TNPs optimize a model’s performance on various tasks by iteratively refining predictions and learning optimal strategies.
Similarly, BONET [184] proposes a generative framework that uses autoregressive models for pretraining optimization on offline datasets. By training a transformer-based model in an unsupervised manner, BONET learns to optimize task performance in real-world settings without direct task-specific supervision. This enables it to adapt to a wide variety of optimization problems, from machine learning to more general generative optimization tasks.
@@ 5 Applications
There are lots of applications of generative models in decision making scenarios. We consider five typical applications including, robot control, autonomous driving, games, structural generation, and optimization.
5.1 Robot Control
Robot control refers to the process of commanding a robot to execute specific tasks or actions [208, 209, 210, 211, 212]. This can be achieved through a variety of methods, including manual control by a human operator, pre-programmed instructions, or autonomous decision-making using sensors and machine learning algorithms.
@@ 5.2 Structural Generation
Generative models are increasingly applied to various graph-based tasks, such as graph generation [219], graph completion [220], and graph classification [221]. In graph generation, these models learn the underlying structures of training graphs and use this knowledge to generate new graphs with similar characteristics. This capability has wide-ranging applications, including molecule design in chemistry, protein interaction modeling in biology, and architecture optimization in engineering.
A notable example of applying generative models in structure generation is through GFlowNets. Jain et al. [222] proposed an active learning framework that integrates epistemic uncertainty estimation with GFlowNets to generate diverse and informative candidate solutions for drug discovery and other problems. This method selects a batch of promising candidates after each learning round. To improve the exploration efficiency, Kim et al. [223] introduced a local search strategy within GFlowNets, targeting high-reward sample spaces while preventing over-exploration. This approach combines forward and backward policies with backtracking and reconstruction techniques, further refining the generated solutions.
Recent works [224, 225, 226] have made notable advancements in integrating reinforcement learning (RL) with natural language processing (NLP). Xin et al. [224] propose Prompt-Based Reinforcement Learning (PRL), which uses historical data and state-reward pairs as prompts to train RL-based recommendation agents. Deng et al. [225] introduce RLPROMPT, an efficient method for discrete prompt optimization that leverages RL to discover optimal prompts for pre-trained language models (LMs). This approach is applicable across different types of LMs and tasks, including both classification and generation. Zhang et al.[226] present TEMPERA, which designs optimal prompts for large LMs in zero-shot or few-shot learning settings by leveraging prior knowledge and enabling flexible prompt edits.
These works illustrate how RL-optimized generative approaches, like prompt-based methods, improve decision-making in language models. By structuring inputs and outputs flexibly, they facilitate context-aware solutions for complex tasks, such as multi-step reasoning and dynamic adaptation, with potential applications beyond NLP in fields like robotics and planning.
@@ 5.3 Games
Games artificial intelligence (AI) is a prominent research area focused on developing AI techniques that achieve human-level performance in gameplay [227]. Games offer intriguing and complex problems for AI agents to solve within safe and controllable environments. They also provide diverse challenges, from strategic planning to real-time decision-making, serving as benchmarks for testing and advancing AI algorithms.
Several generative models have been developed for gameplay. Adversarial Flow Networks (AFlowNets) [170] extend EFlowNets to adversarial environments, specifically for two-player zero-sum games. GATO [179] is a versatile generalist agent capable of playing multiple games. Similarly, the Multi-Game Decision Transformer [180] adapts a single transformer-based model to handle diverse gaming scenarios.
@@ 5.4 Autonomous Driving
Generative models have been used in autonomous driving for a variety of tasks, including driving control [228, 131, 229], image and video processing [229], object detection [230] and scene understanding [231].
Huang et al. [232] introduced a hybrid framework that integrates neural decision-making into the classical modular pipeline using end-to-end imitation learning via reparameterized generative adversarial learning. This approach retains the advantages of the classical pipeline, such as strict adherence to physical and logical constraints, while enabling the framework to learn complex driving decisions directly from data.
Another application of generative model in autonomous driving is in the generation of synthetic data for training deep learning models [233]. Synthetic data helps augment real-world datasets, which are often constrained in size and diversity. These models can produce synthetic images and videos resembling real-world scenarios, thereby enhancing the performance of deep learning models in autonomous driving tasks. For example, TrafficGen [234] acts as an autoregressive generative model based on an encoder-decoder structure. It is trained on fragmented human driving data gathered from real-world environments to create lifelike traffic scenarios. By employing an attention mechanism, TrafficGen encodes the current traffic context. subsequently, it decodes the initial state of the vehicle and generates a long-term trajectory.
Generative models are valuable for processing sensor data, such as LiDAR and camera inputs, in applications like object detection and tracking [235]. These models can generate 3D representations of the environment from sensor data, which serve as a foundation for more effective object detection and tracking.
Generative model can also be used for scene understanding in autonomous driving [96]. For example, it can be used to generate semantic maps of the environment, which can be used for navigation and path planning. Generative model can also be used to generate predictions of future vehicle and pedestrian movements, which can be used for collision avoidance and planning.
PlanCP [236] leverages Conformal Prediction (CP) to quantify the uncertainty in diffusion dynamics models. By utilizing a finite set of exchangeable expert trajectory examples, PlanCP enhances robustness across diverse training scenarios, ensuring more reliable performance in autonomous driving applications.
Overall, generative models hold significant potential for enhancing the performance and safety of autonomous driving systems [237]. They are versatile tools applicable to a range of tasks, including data augmentation and scene understanding, and have shown promising results in both research studies and practical applications.
@@ 5.5 Optimization
We consider three optimization settings: black-box optimization, neural architecture search, and scheduling. These settings represent critical challenges in decision-making tasks, where generative models have demonstrated significant potential.
In the black-box optimization setting, an agent learns to optimize an unknown function through pointwise evaluations [238]. For example, Transformer Neural Process [185] reframes meta learning with uncertainty awareness in the context of sequence modeling, solving it with an autoregressive likelihood-based objective, utilizing a novel transformer-based architecture.
Generative models such as GFlowNets have also been applied to combinatorial optimization. Zhang et al. [239] applied Markov decision processes (MDPs) to various combinatorial problems, training conditional GFlowNets to sample from solution spaces. DIFUSCO [240] introduced a graph-based diffusion framework to solve NP-complete problems by casting them as binary vector optimization tasks and leveraging graph-based denoising diffusion models to create superior methods.
Neural Combinatorial Optimization (NCO) approaches focus on leveraging neural networks to learn feasible solution distributions for combinatorial optimization problems [241, 242]. Autoregressive construction solvers [243, 244, 245, 246, 147] are built upon the success of transformer-based [109] architectures in sequential generation tasks. Non-autoregressive construction solvers [247, 248, 249, 250, 251, 252] have also been proposed to learn high quality solution distributions.
Furthermore, prompt tuning and reinforcement learning (RL)-based prompt optimization methods are practical applications of black-box optimization. Prompt tuning [253, 254, 255, 256] optimizes task-specific embeddings to guide model behavior efficiently, while RL-based methods like RLPROMPT [225] and TEMPERA [226] refine prompt design to improve performance in tasks like classification and few-shot learning. These approaches demonstrate how generative models can navigate vast solution spaces effectively, balancing exploration and exploitation.
In the neural architecture search (NAS) setting. NAS focuses on optimizing neural network designs to maximize task performance. GFlowOut [257] leverages GFlowNets to learn the posterior distribution of dropout masks, providing an innovative method for optimizing network architecture. By utilizing generative models, GFlowOut highlights the potential for improving architecture efficiency while maintaining high performance.
In the scheduling setting, optimization involves designing strategies to allocate resources or order tasks efficiently. GFlowNets have been deployed in compiler scheduling [258], where they control the balance between the diversity and quality of the proposed methods. This application underscores the ability of generative models to manage competing objectives in complex, high-stakes environments, such as industrial and computational settings.
These three optimization settings illustrate the versatility of generative models in tackling diverse and complex challenges, offering promising avenues for future exploration and development.
@@ 6 Perspective: Next Generation Generative Decision Making Models
Despite numerous successful examples, advancing next-generation generative decision-making models remains a critical challenge. In this section, we highlight three pivotal perspectives to enhance the capabilities of generative models: high-performance algorithms, large-scale generalized decision-making Models, and self-evolving and adaptive generative models.
@@ 6.1 High performance Algorithms
The growing integration of generative models in decision-making highlights an increasing demand for high-performance algorithms capable of addressing the complexities of real-world applications. Future research should focus on enhancing their scalability, efficiency, and adaptability to dynamic environments. Normalizing flows and GFlowNets represent two promising directions for advancing generative decision-making algorithms, each with unique strengths and challenges.
Normalizing flows have demonstrated potential in modeling uncertainties and supporting principled decision-making in complex, high-dimensional environments. However, several critical challenges remain. These include designing more effective objective functions tailored to decision-making tasks, improving their efficiency in handling high-dimensional decision spaces, and reducing the computational costs of training. Future work should explore novel architectures and optimization strategies to address these limitations, enabling their broader application in dynamic, real-world scenarios.
Similarly, GFlowNets present a novel probabilistic approach to modeling unnormalized probability distributions, making them particularly effective for generating structured or discrete objects such as graphs. Unlike VAEs and GANs, whose goals are to approximate data distributions or generate realistic samples, GFlowNets are specifically designed to sample $x$ in proportion to a predefined reward function $R(x)$. The objective ensures that the generated samples approximately follow the desired reward distribution.
By leveraging their strengths and addressing current limitations, these frameworks have the potential to reshape generative decision-making, enabling more efficient, scalable, and adaptable systems for complex, dynamic real-world applications.
@@ 6.2 Large-scale Generalized Decision Making Models
As artificial intelligence continues to advance towards large-scale models, developing generative large models for decision-making reasoning becomes increasingly essential. These models have the potential to handle complex reasoning tasks, integrate diverse knowledge, and offer robust solutions in dynamic and uncertain environments.
Additionally, current generative models are often tailored to predefined tasks or scenarios, limiting their generalization capabilities. To address real-world challenges, there is a pressing need for large-scale, generalized generative decision models capable of adapting to diverse and dynamic environments.
For instance, web-scale diffusion models represent a significant step toward achieving this generalization. These models leverage massive datasets and scalable architectures to learn intricate patterns across diverse domains, making them suitable for generating solutions in complex decision-making tasks. Similarly, large language models (LLMs), such as GPT [20] or PaLM [259], showcase the ability to process diverse inputs and perform multi-step reasoning, highlighting their potential as general-purpose decision-making tools. Moreover, vision-language-action (VLA) models push this generalization further by integrating multimodal data, enabling them to interpret, reason, and act across diverse scenarios. These examples underline the transformative potential of generative models in building adaptive, large-scale decision-making frameworks.
Besides, generative models should be capable of producing diverse solutions, enabling them to navigate various possibilities and adapt to the complexities in real-world decision-making scenarios. This diversity is particularly important in handling ambiguous or multi-objective tasks, where a single solution may not suffice. By generating varied yet high-quality outputs, these models can better account for uncertainties, provide alternative strategies, and support robust decision-making processes. Such capabilities are critical for applications in dynamic and multi-faceted domains, ranging from autonomous systems to large-scale optimization tasks.
@@ 6.3 Self-Evolving and Adaptive Generative Models
The third perspective emphasizes the importance of self-evolving capabilities in generative models to address dynamic and complex decision-making environments. While current generative models are often built with static assumptions and trained on fixed datasets, real-world decision-making scenarios require models that can adapt continuously, learn in real time, and transfer knowledge across different domains.
A key component of adaptive generative models is continual learning, which allows them to evolve as new data becomes available. This ability allows them to update their knowledge without forgetting previously learned information, a challenge known as catastrophic forgetting. By leveraging incremental training, memory replay, and regularization techniques, these models can respond to evolving data distributions and changing environments, ensuring their decision-making remains accurate and relevant. This adaptability is crucial for applications where data patterns shift frequently and models need to stay up-to-date without retraining from scratch.
Furthermore, the integration of real-time feedback mechanisms allows generative models to quickly adjust their decision-making strategies based on immediate outcomes or user interactions. Real-time performance monitoring can trigger dynamic updates, enabling the model to refine its approach as it receives feedback from the environment. This is essential for tasks that require quick reactions, such as autonomous navigation or adaptive recommendation systems, where rapid changes in input or context need to be addressed effectively to maintain high performance.
Cross-domain adaptability is another vital aspect, as real-world decision-making often spans multiple domains with distinct challenges and data characteristics. Generative models that can transfer knowledge learned from one domain to another can enhance their versatility and expand their applicability. Approaches like meta-learning and transfer learning enable models to recognize and apply shared structures and features across different tasks, making them capable of tackling new and unseen problems with minimal additional training. This cross-domain capability ensures that models are not limited to a narrow set of conditions but are instead able to operate effectively in diverse and multi-faceted environments.
Incorporating adaptive features like continual learning, real-time feedback, and cross-domain adaptability allows generative models to better align with real-world challenges. This approach paves the way for more robust, versatile, and future-proof decision-making systems that can thrive in complex and ever-changing environments, extending the potential of generative models far beyond their current scope and applications.

### v3

@@ 4.3.7 Autoregressive Models as Optimizer
Autoregressive models are effective optimizers due to their sequential structure, which captures dependencies between data points, allowing for iterative solution refinement. This ability helps them handle uncertainty and improve optimization performance across various tasks. Their sequential nature enables them to generate or adjust solutions gradually, making them suitable for complex, high-dimensional solution spaces.
For example, Transformer Neural Processes (TNPs) [172] use a transformer-based autoregressive model within a meta-learning framework to optimize performance in uncertain tasks with variable inputs. By combining transformers’ strengths in sequence modeling with Neural Processes for uncertainty estimation, TNPs refine predictions and learn optimal strategies iteratively. (see in Fig. 14)
Similarly, BONET [171] introduces a generative framework using autoregressive models to pretrain optimization on offline datasets. The transformer-based model learns to optimize task performance in real-world settings without task-specific supervision, enabling adaptation to a wide range of optimization problems.
@@ 5 Applications
There are lots of applications of generative models in decision making scenarios. We consider five typical applications including: robot control, autonomous driving, games, structural generation, and optimization.
5.1 Robotic Control
Robotic control refers to the process of commanding a robot to execute specific tasks or actions [191, 192, 193, 194, 195]. This can be achieved through a variety of methods, including manual control by a human operator, pre-programmed instructions, or autonomous decision-making using sensors and machine learning algorithms.
@@ 5.2 Structural Generation
Generative models are widely used in graph-based tasks, such as graph generation [202], completion [203], and classification [204]. These models learn the structure of training graphs and generate new ones with similar characteristics, which has applications in fields like molecule design, protein interaction modeling, and architecture optimization.
One key example is the use of Generative Flow Networks (GFlowNets) for structure generation. Jain et al. [205] introduced an active learning framework with GFlowNets to generate diverse candidate solutions for drug discovery. Kim et al. [206] enhanced exploration by incorporating a local search strategy, combining forward and backward policies to improve solution quality.
Recent work integrating reinforcement learning (RL) with natural language processing (NLP) has further advanced structural generation for decision-making. Xin et al. [207] proposes Prompt-Based Reinforcement Learning (PRL) to optimize decision-making tasks by using state-reward pairs as prompts. Deng et al. [208] introduces RLPROMPT, an RL-based method for prompt optimization in pre-trained language models. Zhang et al. [209] extends this with TEMPERA, designing optimal prompts for zero-shot and few-shot learning.
These advances demonstrate how RL-enhanced generative models improve decision-making in NLP, and their potential extends to areas like robotics and planning.
@@ 5.3 Games
Games AI is a key research area aimed at developing AI systems that can perform at human-level across various gameplay types [210]. Games offer complex, controlled environments with challenges ranging from strategic planning to real-time decision-making, making them ideal benchmarks for AI development.
For single-player games, generative models have been increasingly applied [211]. The Multi-Game Decision Transformer [166] utilizes a transformer-based model to efficiently handle various game scenarios, highlighting the flexibility of generative approaches for diverse gaming tasks. Another generalist agent, GATO [165], is capable of playing a variety of games, demonstrating the potential for generative models to adapt across multiple player environments.
In multiplayer games [212, 213, 157, 213], Adversarial Flow Network (AFlowNet) [157] extends EFlowNets to adversarial two-player zero-sum games. On the other hand, Generative Agents [213] use generative models to simulate human-like behaviors and interactions among multiple agents in dynamic virtual environments, enhancing the complexity and depth of multiplayer experiences.
@@ 5.4 Autonomous Driving
Generative models have shown great potential in autonomous driving, particularly in driving control [214, 118, 215], object detection [216], and scene understanding [217]. They are especially effective in driving decision-making, generating complex policies from data. Huang et al. [218] proposed a hybrid framework that combines neural decision-making with classical pipelines, enabling dynamic decisions while adhering to physical constraints.
A major challenge in autonomous driving is the lack of diverse training data. Generative models address this by producing synthetic data to simulate real-world traffic scenarios. For example, TrafficGen [219] uses an autoregressive model to generate long-term driving trajectories, helping overcome data scarcity, especially in edge cases.
Generative models also improve sensor data processing, enhancing object detection and tracking from LiDAR and camera inputs [220], by creating 3D environment representations for more accurate real-time decisions.
In scene understanding and path planning, these models generate semantic maps and predict object movements for safe navigation. PlanCP [221] enhances this by using Conformal Prediction (CP) to quantify uncertainty and assess policy robustness across scenarios.
Despite challenges like ensuring data quality and adapting to dynamic real-world conditions, generative models are essential for autonomous driving systems, offering scalable, safe, and robust solutions [222].
@@ 5.5 Optimization
Generative models show significant potential in three optimization settings: black-box optimization, neural architecture search (NAS), and scheduling optimization.
Neural Combinatorial Optimization (NCO) leverages neural networks to learn solution distributions for combinatorial problems [223, 224]. This includes autoregressive [225, 226, 227] and non-autoregressive solvers [228, 229, 230] based on transformer architectures. GFlowNets [231] and methods like DIFUSCO [232] apply graph-based diffusion models to solve NP-complete problems.
In black-box optimization, generative models optimize task-specific embeddings through prompt tuning [233, 234, 235, 236], improving model performance. RL-based methods like RLPROMPT [208] and TEMPERA [209] further refine prompt design, enhancing tasks such as classification and few-shot learning.
In NAS, generative models optimize neural network designs to maximize performance. GFlowOut [237] uses GFlowNets to optimize architectures by learning posterior distributions of dropout masks. GFlowNets are also applied in scheduling optimization, such as compiler scheduling [238], balancing the diversity and quality of proposed solutions in complex environments.
These optimization settings highlight the versatility of generative models in tackling diverse challenges, offering exciting prospects for further exploration and development.
@@ 6 Perspective: Next Generation Generative Decision Making Models
Despite numerous successes, advancing next-generation generative decision-making models remains a critical challenge. In this section, we highlight three pivotal perspectives to enhance the generative models: high-performance algorithms, large-scale generalized decision-making models, and self-evolving and adaptive generative models.
@@ 6.1 High performance algorithms
The growing use of generative models in decision-making highlights an increasing demand for high-performance algorithms capable of addressing the complexities of real-world applications. Future research should focus on improving scalability, efficiency, and adaptability to dynamic environments. Two promising directions are Normalizing Flows and GFlowNets, each with unique strengths and challenges. Normalizing flows excel in modeling uncertainties and supporting decision-making in high-dimensional environments. However, challenges remain, such as designing effective objective functions, improving efficiency in high-dimensional spaces, and reducing training costs. Future work should explore novel architectures and optimization strategies to overcome these limitations.
GFlowNets offer a probabilistic approach to modeling unnormalized distributions, ideal for generating structured objects like graphs. Unlike VAEs and GANs, which approximate data distributions, GFlowNets sample according to a predefined reward function, ensuring the generated samples follow the desired reward distribution.
By addressing these challenges, both frameworks can advance generative decision-making, enabling more efficient, scalable, and adaptable systems for real-world applications.
@@ 6.2 Large-scale Generalized Decision Making Models
As artificial intelligence advances, developing large generative models for decision-making becomes essential to handle complex reasoning, integrate diverse knowledge, and offer robust solutions in dynamic, uncertain environments.
Current generative models are often task-specific, which limits their ability to generalize. To address this, there is a need for large-scale, generalized generative decision models that can adapt to diverse, real-world environments.
For instance, web-scale diffusion models for robotics [239, 185, 240] represent a significant step toward achieving this generalization. These models leverage massive datasets and scalable architectures to learn intricate patterns across diverse domains, making them suitable for generating solutions in complex decision-making tasks. Moreover, vision-language-action (VLA) models [241, 242, 243, 244] push this generalization further by integrating multimodal data, enabling them to interpret, reason, and act across diverse scenarios. These examples underline the transformative potential of generative models in building adaptive, large-scale decision-making frameworks.
Besides, generative models should be capable of producing diverse solutions to handle ambiguous or multi-objective tasks. This diversity allows them to account for uncertainties, offer alternative strategies, and support robust decision-making, critical for applications in dynamic domains like autonomous systems and optimization.
@@ 6.3 Self-Evolving and Adaptive Generative Models
The third perspective highlights the importance of self-evolving capabilities in generative models to address dynamic decision-making environments. While current models are often static, real-world scenarios require systems that can continuously adapt, learn in real time, and transfer knowledge across domains.
A crucial aspect of adaptive models is continual learning, enabling them to evolve as new data arrives without forgetting previously learned information. Techniques like incremental training, memory replay, and regularization help mitigate catastrophic forgetting, allowing models to stay accurate and relevant as data distributions change. This adaptability is essential for applications where data patterns shift frequently.
Real-time feedback further enhances adaptability, enabling models to adjust strategies based on immediate outcomes or user interactions. This feature is critical for tasks like autonomous navigation or adaptive recommendation systems, where rapid responses are necessary to maintain high performance.
Cross-domain adaptability also plays a key role, as decision-making often spans multiple areas with distinct challenges. Models that can transfer knowledge across domains can tackle new problems with minimal additional training. Methods like transfer learning and meta-learning help models identify and apply shared structures, making them versatile across various tasks.
Incorporating continual learning, real-time feedback, and cross-domain adaptability will enable generative models to better align with real-world challenges, paving the way for more robust, versatile decision-making systems that can thrive in complex, evolving environments.